Compute for the pH of an acetate buffer produced by mixing 500 ml of a 0.5 M NaHCO3

solution and 300 ml of 0.25 M H2CO3 solution.


pka = 6.1


1
Expert's answer
2021-03-22T06:25:36-0400

Mmol of Acid = 300×0.25 = 75 mmol

Mmol of Salt = 500×0.50 = 250 mmol

It is Acidic buffer

"pH = pKa + log(S\/A)"

pH = 6.1 + log(250/75)

= [6.62]